OpenStudy (anonymous):

y^(5)(x) + y'(x) = 0 find the general solution

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

ofcourse, i meant not the identically 0 function that is a solution to all homogeneous diff eqs...

OpenStudy (nikvist):

\[y^5+y'=0\]\[1+y^{-5}y'=0\]\[-\frac{1}{4}(y^{-4})'=-1\]\[(y^{-4})'=4\]\[y^{-4}=4x+C\]\[y^{4}=\frac{1}{4x+C}\]